OpenStudy (anonymous):

The product of (a + b)(a - b) is a perfect square trinomial. Sometimes Always Never

OpenStudy (anonymous):

@adam.bloedel

OpenStudy (solomonzelman):

(a+b)(a-b)=a²-b²

OpenStudy (juanpablojr):

Seems like a binomial.

OpenStudy (anonymous):

sorry i dont really get that this is really hard for me and i know that is a binomial

OpenStudy (juanpablojr):

Since its a binomial how can it be a Trinomial.

OpenStudy (anonymous):

thank you so so so much :)

OpenStudy (anonymous):

can you help on one more pwease?

OpenStudy (juanpablojr):

Welcome :) and ill see if I can

OpenStudy (anonymous):

okay here it is Find the product of (x + 5)(x - 5). x2 - 10x + 25 x2 + 10x + 25 x2 - 25 x2 + 25 im guessing its the last one

OpenStudy (solomonzelman):

(a+b)(a+b) = (a+b)² = a²+2ab+b²

OpenStudy (juanpablojr):

I just wrote that lol

OpenStudy (anonymous):

what do i do since it is +5 and -5

OpenStudy (anonymous):

its the first one right?

OpenStudy (juanpablojr):

you know what FOIL is?

OpenStudy (anonymous):

um sort of

OpenStudy (juanpablojr):

First Outside Inside Last Multiply the first terms in each bracket Then Multiply the outside terms, Inside then last.

OpenStudy (anonymous):

i still think it is the first one

OpenStudy (juanpablojr):

x2 is x^2 right as in x squared?

OpenStudy (anonymous):

yeah sorry about that

OpenStudy (juanpablojr):

(x - 5)(x + 5) so x(x) + x(5) - 5(x) - 5(5) by foil

OpenStudy (juanpablojr):

Then you sub it to x² + 5x - 5x + 25 then you combine like terms to get your answer.

OpenStudy (anonymous):

i got the last one thank you so much love you are a lifesaver :)

OpenStudy (juanpablojr):

NOOO you should \[X ^{2}-25\]

OpenStudy (juanpablojr):

wait I might of made a mistake did u get it right?

OpenStudy (anonymous):

uh oh.... i already submitted it and got an 80 but thats fine thanks

OpenStudy (juanpablojr):

Oh just remember foil, substitute and combine the like terms and you should be fine. :)